Wednesday, April 13, 2011

Day 15, 04/13/2011 -- Something is rotten in the state of Texas

Rotten is perhaps too strong a word. Basically, when you do experimental physics, things often don't work. It's part of the simultaneous charm and frustration of doing new things.

In this case, we are having trouble with the "grippers", a set of three motor-controlled jaws that hold the half-wave-plate in place as the cryostat cools down. Despite repeated testing in another cryostat and Jeff's careful assembly, we have yet to succeed in opening the grippers and releasing the HWP. We still have a few things to try tomorrow, so hope is far from lost. Even if the HWP isn't free to rotate, we still have plenty of other tests that need to be done, both without and eventually with the gondola, so we are forging ahead.

While the UMN crew tackled the gripper problem, Franky kept himself busy with continuing to assemble and populate the BROs. At this point, we're just waiting for the detectors to be cold enough that we can start connecting cables and doing the baseline checkout tests.

After Jeff and I went in at 4:30 AM for a helium transfer, the boiloff rate is now slowing down and will likely last through the night -- a good sign that we are close to being able to run the fridges and detectors.

Pictures:
https://picasaweb.google.com/asad137/EBEXInPalestine20110413

No comments:

Post a Comment